OpenStudy (anonymous):

What is the value of the expression 11 + (fraction 1 over 2)4 ⋅ 48? direct answer pls

OpenStudy (alexandervonhumboldt2):

we cannot give out direct answers it is against the rules. we give only explanations

OpenStudy (anonymous):

ok can u help me

OpenStudy (alexandervonhumboldt2):

ok so first we do multiplying 11+1/2*4*48 1/2*4 is the same as 4/2=2 so we have 11+2*48. now again multiply 2 by 48. what do we get when we 2*48=?

OpenStudy (alexandervonhumboldt2):

what is 2*48 @galaxystare ?

OpenStudy (anonymous):

96

OpenStudy (alexandervonhumboldt2):

yeah so we have 11+98=???

OpenStudy (anonymous):

109

OpenStudy (alexandervonhumboldt2):

yes so that is your answer

OpenStudy (anonymous):

these r the answers 12 14 17 19

OpenStudy (alexandervonhumboldt2):

well then you incorrectly typed your question check it again

OpenStudy (anonymous):

no i copied and pasted

OpenStudy (alexandervonhumboldt2):

well then you incorrectly typed your question check it again

OpenStudy (alexandervonhumboldt2):

@ganeshie8 @hartnn @wio @freckles @Jhannybean